$SPAD/src/input as-eg5.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
)co pdecomp0.as
(u,v):UP(x,FRAC INT)
u:=x**2+x+2
v:=x**3+2*x**2+x+1
compose(u,v)
decompose %

———-

References^{[1] nothing}

4